OpenStudy (anonymous):

Use this coordinate ,find the other coordinate that makes the ordered pair a solution of the given equation: 3x+4y=15,(_,7/2)

OpenStudy (anonymous):

So you have the equation 3x+4y=15. You are given the value of y in (x,7/2). So you are solving for the x coordinate that satisfies the equation when y=7/2. First solve the equation for x: x=(15-4y)/3. Now plug in the y coordinate you were given: x=(15-4(7/2))/3=1/3. So the x coordinate is 1/3. In other words the point (1/3,7/2) is on the line 3x+4y=15.